OpenStudy (anonymous):

3.25/5.789=????? answer to 3 decimal place

OpenStudy (ookawaiioo):

use a calculator LOL X.XXX it should look like that but with numbers

OpenStudy (anonymous):

\[\frac{ 3.25 }{ 5.789 }=\frac{ 3.250 }{5.789 }=\frac{ 3250 }{5789 }\] now you can divide (log division ) or use calculator.